OpenStudy (anonymous):

The perimeter of a rectangular field is 358 yards. If the length of the field is 96 yards, what is its width?

OpenStudy (jango_in_dtown):

2(length+width)=perimeter

OpenStudy (jango_in_dtown):

put the value of length and perimeter from the given question

OpenStudy (anonymous):

Simple, perimeter=2L+2W. If the length is 96 yards, put that into the equation. p=2(96)+2w 358=192+2w -192 166=2w divide by two. w=83 yards.

OpenStudy (jango_in_dtown):

so 2(96+width)=358

OpenStudy (jango_in_dtown):

just solve for width